# Federated Graph Condensation with Information Bottleneck Principles

### v3

@@ 4.2 Overall Performance
Performance on MIA. We replace IB with other defense baselines to evaluate the effects of defending MIA. From Table 2, we observe although the baselines can achieve comparable results with FedGC on the utility of condensed graphs, they suffer a dramatic performance decline in defending MIA. By contrast, FedGC can simultaneously achieve high utility and preserve membership privacy. Since the condensed graph is public, model-based defenses (Reg and LDP) can only affect the model gradients matching stage, which has little intervention on the properties entailed from original graphs, leading to a high success of MIA. In terms of this, GCN+PL directly modifies the original data (adding pseudo labels), thus obtaining better defense. However, the pseudo labels inherit the knowledge of training labels, making MIA still effective. Leveraging IB principles, FedGC transforms original node features into a new space, preventing strong attackers from concluding memberships.
@@ 4.4 Ablation Study
Ablation study on MIA. The ablation study on MIA is depicted in Figure 5. Results on other datasets are presented in the Appendix. We remove the -ft because it’s conducted in the test phase and has no impact on MIA. We can see that the local graph transformation with IB has a significant defensive ability to MIA (up to 75% MIA performance decay on Citeseer). As discussed before, the optimization of the IB objective in Eq. (5) will minimize inference cost gain, making the MIA less effective. It can also be seen that one-step communication and self-training defend MIA to some degree. One-step communication can be seen as a feature augmentation and self-training aims to augment labels. These augmentations make the condensed graph fit to augmented data rather than original data thus alleviating the over-fitting, leading to a performance decay of MIA.
@@ Appendix E Implementation Details
The condensation ratio $r$ is defined as the proportion of condensed nodes to original nodes (Jin et al. 2022b). Following (Jin et al. 2022b; Zheng et al. 2023) ,for small dataset Cora and Citeseer, we choose $r$ as {1.3%, 2.6%, 5.2%} and {0.9%, 1.8%, 3.6%} respectively. For large-scale datasets Ogbn-arxiv, Flickr and Reddit, we choose $r$ as {0.05%, 0.25%, 0.5%}, {0.1%, 0.5%, 1%} and {0.05%, 0.1%, 0.2%} respectively. For all the baselines, the node features are randomly initialized and the hidden dimension is set to 256. For FedGC, we utilize a 2-layer GCN (Kipf and Welling 2017) in self-training phrase and test phrase. For the condensation methods, we use a 2-layer GCN as the backbone (3-layer SGC (Wu et al. 2019) for Ogbn-arxiv). We use Mean Square Error (MSE) as the distance function of gradient matching for all dataset, except for using the distance from (Jin et al. 2022b) in Ogbn-arxiv:
@@
For MIA, following (He et al. 2021), we assume that the attacker possesses a subgraph of the dataset with a small set of labels to form the shadow dataset. The nodes used for training and testing shadow models have no overlaps with original train and test nodes. They are all randomly sampled with the same size. The attack model is a 3-layer MLP and the shadow and target model are 2-layer GCN. For all the We tune other hyper-parameters based on valid datasets to report the best performance.
@@ F.1 Generalizability of condensed graphs on more datasets
Table 5 presents the generalizability of condensed graphs in Flickr and Reddit datasets. We can see that FedGC outperforms GCOND and SFGC w.r.t all the GNN architectures and even the simple MLP. The improvements is significant, demonstrating the well generalizability of condensed graph by FedGC, especially in large-scale graph datasets.
@@ Appendix G Related work
Dataset condensation. Due to the prohibitive costs of training models with large-scale datasets, dataset condensation (DC) has emerged to synthesize a smaller dataset by distilling the knowledge of a given large dataset. Subsequently, the condensed dataset can be served as a substitute of the large dataset to conduct downstream tasks, such as hyper-parameter tuning and neural architecture search (NAS). Typically, dataset distillation (DD) (Wang et al. 2018) firstly achieves this by a learning-to-learn framework with bi-level optimization. To alleviate the high computational overhead brought by unrolled computation graph, DC (Zhao, Mopuri, and Bilen 2021; Zhao and Bilen 2021) is proposed to match the gradients between the original graph and the condensed graph. Similar idea is also adopted to match the distributions (Zhao and Bilen 2023) and training trajectories (Cazenavette et al. 2022). Extending to the graph data, GCOND (Jin et al. 2022b) condenses a large graph into a smaller one for node classifications based on gradient matching scheme. DosCond (Jin et al. 2022a) further proposes a one-step gradient matching algorithm to avoid nested optimization of GCOND. It is also compatible with graph classification tasks. Other works attempt to match the distributions (Liu et al. 2022) and training trajectories (Zheng et al. 2023) of graph data and achieve remarkable results. Despite their success, they focus on the graph condensation under centralized data storage, regardless of restrained conditions with distributed data or locally privacy-preserving requirements.
@@
Federated graph learning. Federated learning (FL) (McMahan et al. 2017; Yang, Fan, and Yu 2020) has emerged to endow the ability of collaboratively learning a global model with distributed data while protecting data privacy. Federated graph learning (FGL) extends the FL to the graph domain and can be divided into inter-graph and intra-graph learning based on the data distributed type (Zhang et al. 2021a). In inter-graph learning, each client owns multiple graphs and aims to train a global model with other clients’ cooperation for graph-level tasks (Xie et al. 2021), such as molecular property prediction (Zhu, Luo, and White 2022). In contrast, intra-graph learning assumes each client only owns a subgraph of the whole graph, which therefore mainly focuses on recovering missing cross-client edges and neighbors. BDS-GCN (Wan et al. 2022) proposes sampling cross-client neighbors but violates other clients’ privacy. Instead, FedSage+ (Zhang et al. 2021b) generates missing neighbors locally by training linear generators. However, it relies on communications between clients in each federated training round. To reduce communication cost, FedGCN (Yao et al. 2023) exchanges neighbor aggregations through homomorphic encryption in one pre-training step, largely speeding up federated training process and also protecting local privacy. Some other intra-graph learning works dedicate to tackling the heterogeneity issues among clients (Baek et al. 2023; Tan et al. 2023). They model the similarities or share common knowledge between clients to alleviate the performance degradation caused by data heterogeneity. All these existing federated graph learning methods require burdensome federated training process, which even need to repeat multiple times in some downstream tasks, such as hyper-parameter tuning and neural architecture search. In light of this limitation, in this paper we focus on intra-graph learning and study federated graph condensation to obtain a global small graph for facilitating downstream tasks.
@@
Membership inference attack. Membership inference attack (MIA)(Salem et al. 2019; Nasr, Shokri, and Houmansadr 2018; Carlini et al. 2019; Chen et al. 2023; Zhang et al. 2021c) is a privacy attack enabling an attacker to illicitly obtain private or sensitive data from a learning model, by discerning the use of a specific sample as training data. Shadow training is generally used to train the attack model for MIA (Shokri et al. 2017; Salem et al. 2019), in the case that the training data with some background knowledge, such as full output knowledge, partial output knowledge and label-only knowledge, is queried by the attacker to be the target posterior, the attacker trains multiple shadow models to simulate the target model and uses multiple attack models to attack. There may be differences in the setting of adversarial shadowing model learning attacks, the target victim model access may be black-box or white-box for the attacker(Li, Yi, and Zhang 2018). In a white-box setting, the adversary has complete access to the target victim model, leveraging its entire structure to create adversarial examples and attack models, typically resulting in a near 100% attack success rate. When the adversary has black-box access, which is much more practical, it can only query the target model without knowledge of the model’s structure or parameters (e.g., weights, gradients) or underlying assumptions regarding the model architecture. Several works leverage data augmentations (Kahla et al. 2022) and confident scores (Salem et al. 2019) from the victim model’s output to readily reconstruct local sensitive data, posing a threat to vulnerable learning models on the server. Recently, researches (Olatunji, Nejdl, and Khosla 2021; He et al. 2021) have indicated that GNN are vulnerable to MIA, and the graph properties such as connectivity information encoded by the GNN inside the model structure makes it more vulnerable to such inference attacks. Their attacks conducted in a restricted scenario, using a target node’s subgraph to query the target model to obtain the input to their attack model.
@@
Graph Neural Network. Graph Neural Networks (GNNs) have emerged as a powerful tool for modeling structure data and have been applied in various fields such as recommendation systems (Fan et al. 2019), bioinformatics (Fout et al. 2017), and cybersecurity (Qin, Xiao, and Liao 2023). As a most popular GNNs, GCN (Kipf and Welling 2017) simplifies the convolution process on graphs by using an efficient layer-wise propagation rule. SGC (Wu et al. 2019) removes the nonlinear activation function in GCN to speed up the training process and also keep the competitive performance. GAT (Velickovic et al. 2018) further incorporates attention mechanisms to weigh the influence of nodes during the feature aggregation phase. GraphSage (Hamilton, Ying, and Leskovec 2017) enables inductive learning on graphs by sampling and aggregating features from local neighborhood. To solve the over-smoothing problem of GNNs, APPNP (Klicpera, Bojchevski, and Günnemann 2019) is proposed decoupling the feature transformation from the propagation with an adaptation of personalized PageRank. Besides above mentioned spatial GNNs, there are some GNNs leverage the spectrum of graph Laplacian to perform convolutions in the spectral domain (Bo et al. 2023, 2021). For example, ChebyNet (Defferrard, Bresson, and Vandergheynst 2016) is an implementation method of CNNs on graph with fast localized spectral filtering. In this paper we use GNNs as backbone model for graph condensation and test the performance of condensed graph under different GNN architectures.

### v4

@@ 4.2 Overall Performance
Performance on MIA. We replace IB with other defense baselines to evaluate the effects of defending MIA. From Table 2, we observe although the baselines can achieve comparable results with FedGC on the utility of condensed graphs, they suffer a dramatic performance decline in defending MIA. By contrast, FedGC can simultaneously achieve high utility and preserve membership privacy. Since the condensed graph is public, model-based defenses (Reg and LDP) can only affect the model gradients matching stage, which has little intervention on the properties entailed from original graphs, leading to a high success of MIA. In terms of this, PL directly modifies the original data (adding pseudo labels), thus obtaining better defense. However, the pseudo labels inherit the knowledge of training labels, making MIA still effective. Leveraging IB principles, FedGC transforms original node features into a new space, preventing strong attackers from concluding memberships.
@@ 4.4 Ablation Study
Ablation study on MIA. The ablation study on MIA is depicted in Figure 5. Results on other datasets are presented in the Appendix. We remove the -ft because it’s conducted in the test phase and has no impact on MIA. We can see that the local graph transformation with IB has a significant defensive ability to MIA (up to 40% MIA performance decay on Citeseer). As discussed before, the optimization of the IB objective in Eq. (5) will minimize inference cost gain, making the MIA less effective. It can also be seen that one-step communication and self-training defend MIA to some degree. One-step communication can be seen as a feature augmentation and self-training aims to augment labels. These augmentations make the condensed graph fit to augmented data rather than original data thus alleviating the over-fitting, leading to a performance decay of MIA.
@@ 5 Conclusion
6 Acknowledgments
This work is supported in part by the National Key Research and Development Program of China (2023YFF0725103), the National Natural Science Foundation of China (U22B2038, 62192784), the JSPS KAKENHI JP23K24851, the JST PRESTO JPMJPR23P5, and the JST CREST JPMJCR21M2.
@@ Appendix E Implementation Details
The condensation ratio $r$ is defined as the proportion of condensed nodes to original nodes (Jin et al. 2022b). Following (Jin et al. 2022b; Zheng et al. 2023) ,for small dataset Cora and Citeseer, we choose $r$ as {1.3%, 2.6%, 5.2%} and {0.9%, 1.8%, 3.6%} respectively. For large-scale datasets Ogbn-arxiv, Flickr, and Reddit, we choose $r$ as {0.05%, 0.25%, 0.5%}, {0.1%, 0.5%, 1%} and {0.05%, 0.1%, 0.2%} respectively. For all the baselines, the node features are randomly initialized and the hidden dimension is set to 256. For FedGC, we utilize a 2-layer GCN (Kipf and Welling 2017) in the self-training phrase and test phrase. For the condensation methods, we use a 2-layer GCN as the backbone (3-layer SGC (Wu et al. 2019) for Ogbn-arxiv). We use Mean Square Error (MSE) as the distance function of gradient matching for all datasets, except for using the distance from (Jin et al. 2022b) in Ogbn-arxiv:
@@
For MIA, following (He et al. 2021), we assume that the attacker possesses a subgraph of the dataset with a small set of labels to form the shadow dataset. The nodes used for training and testing shadow models have no overlaps with the original train and test nodes. They are all randomly sampled with the same size. The attack model is a 3-layer MLP and the shadow and target model are 2-layer GCN. For all the baselines, we tune other hyper-parameters based on valid datasets to report the best performance.
@@ F.1 Generalizability of condensed graphs on more datasets
Table 5 presents the generalizability of condensed graphs in Flickr and Reddit datasets. We can see that FedGC outperforms GCOND and SFGC w.r.t all the GNN architectures and even the simple MLP. The improvements are significant, demonstrating the good generalizability of the condensed graph by FedGC, especially in large-scale graph datasets.
@@ Appendix G Related work
Dataset condensation. Due to the prohibitive costs of training models with large-scale datasets, dataset condensation (DC) has emerged to synthesize a smaller dataset by distilling the knowledge of a given large dataset. Subsequently, the condensed dataset can serve as a substitute for the large dataset to conduct downstream tasks, such as hyper-parameter tuning and neural architecture search (NAS). Typically, dataset distillation (DD) (Wang et al. 2018) firstly achieves this by a learning-to-learn framework with bi-level optimization. To alleviate the high computational overhead brought by the unrolled computation graph, DC (Zhao, Mopuri, and Bilen 2021; Zhao and Bilen 2021) is proposed to match the gradients between the original graph and the condensed graph. Similar idea is also adopted to match the distributions (Zhao and Bilen 2023) and training trajectories (Cazenavette et al. 2022). Extending to the graph data, GCOND (Jin et al. 2022b) condenses a large graph into a smaller one for node classifications based on gradient matching scheme. DosCond (Jin et al. 2022a) further proposes a one-step gradient matching algorithm to avoid nested optimization of GCOND. It is also compatible with graph classification tasks. Other works attempt to match the distributions (Liu et al. 2022) and training trajectories (Zheng et al. 2023) of graph data and achieve remarkable results. Despite their success, they focus on the graph condensation under centralized data storage, regardless of restrained conditions with distributed data or local privacy-preserving requirements.
@@
Federated graph learning. Federated learning (FL) (McMahan et al. 2017; Yang, Fan, and Yu 2020) has emerged to endow the ability of collaboratively learning a global model with distributed data while protecting data privacy. Federated graph learning (FGL) extends the FL to the graph domain and can be divided into inter-graph and intra-graph learning based on the data distributed type (Zhang et al. 2021a). In inter-graph learning, each client owns multiple graphs and aims to train a global model with other clients’ cooperation for graph-level tasks (Xie et al. 2021), such as molecular property prediction (Zhu, Luo, and White 2022). In contrast, intra-graph learning assumes each client only owns a subgraph of the whole graph, which therefore mainly focuses on recovering missing cross-client edges and neighbors. BDS-GCN (Wan et al. 2022) proposes sampling cross-client neighbors but violates other clients’ privacy. Instead, FedSage+ (Zhang et al. 2021b) generates missing neighbors locally by training linear generators. However, it relies on communications between clients in each federated training round. To reduce communication costs, FedGCN (Yao et al. 2023) exchanges neighbor aggregations through homomorphic encryption in one pre-training step, largely speeding up the federated training process and also protecting local privacy. Some other intra-graph learning works are dedicated to tackling the heterogeneity issues among clients (Baek et al. 2023; Tan et al. 2023). They model the similarities or share common knowledge between clients to alleviate the performance degradation caused by data heterogeneity. All these existing federated graph learning methods require burdensome federated training processes, which even need to be repeated multiple times in some downstream tasks, such as hyper-parameter tuning and neural architecture search. In light of this limitation, we focus on intra-graph learning and study federated graph condensation to obtain a global small graph for facilitating downstream tasks.
@@
Membership inference attack. Membership inference attack (MIA)(Salem et al. 2019; Nasr, Shokri, and Houmansadr 2018; Carlini et al. 2019; Chen et al. 2023; Zhang et al. 2021c) is a privacy attack enabling an attacker to illicitly obtain private or sensitive data from a learning model, by discerning the use of a specific sample as training data. Shadow training is generally used to train the attack model for MIA (Shokri et al. 2017; Salem et al. 2019), in the case that the training data with some background knowledge, such as full output knowledge, partial output knowledge, and label-only knowledge, is queried by the attacker to be the target posterior, the attacker trains multiple shadow models to simulate the target model and uses multiple attack models to attack. There may be differences in the setting of adversarial shadowing model learning attacks, the target victim model access may be black-box or white-box for the attacker(Li, Yi, and Zhang 2018). In a white-box setting, the adversary has complete access to the target victim model, leveraging its entire structure to create adversarial examples and attack models, typically resulting in a near 100% attack success rate. When the adversary has black-box access, which is much more practical, it can only query the target model without knowledge of the model’s structure or parameters (e.g., weights, gradients) or underlying assumptions regarding the model architecture. Several works leverage data augmentations (Kahla et al. 2022) and confident scores (Salem et al. 2019) from the victim model’s output to readily reconstruct local sensitive data, posing a threat to vulnerable learning models on the server. Recently, researches (Olatunji, Nejdl, and Khosla 2021; He et al. 2021) have indicated that GNN is vulnerable to MIA, and the graph properties such as connectivity information encoded by the GNN inside the model structure makes it more vulnerable to such inference attacks. Their attacks are conducted in a restricted scenario, using a target node’s subgraph to query the target model to obtain the input to their attack model.
@@
Graph Neural Network. Graph Neural Networks (GNNs) (Wu et al. 2021; Zhang et al. 2024; Xu et al. 2019) have emerged as a powerful tool for modeling structure data and have been applied in various fields such as recommendation systems (Fan et al. 2019), bioinformatics (Fout et al. 2017), and cybersecurity (Qin, Xiao, and Liao 2023). As a most popular GNN, GCN (Kipf and Welling 2017) simplifies the convolution process on graphs by using an efficient layer-wise propagation rule. SGC (Wu et al. 2019) removes the nonlinear activation function in GCN to speed up the training process and also keep the competitive performance. GAT (Velickovic et al. 2018) further incorporates attention mechanisms to weigh the influence of nodes during the feature aggregation phase. GraphSage (Hamilton, Ying, and Leskovec 2017) enables inductive learning on graphs by sampling and aggregating features from local neighborhoods. To solve the over-smoothing problem of GNNs, APPNP (Klicpera, Bojchevski, and Günnemann 2019) is proposed to decouple the feature transformation from the propagation with an adaptation of personalized PageRank. Besides above mentioned spatial GNNs, there are some GNNs that leverage the spectrum of graph Laplacian to perform convolutions in the spectral domain (Bo et al. 2023, 2021). For example, ChebyNet (Defferrard, Bresson, and Vandergheynst 2016) is an implementation method of CNNs on graphs with fast localized spectral filtering. In this paper, we use GNNs as a backbone model for graph condensation and test the performance of the condensed graph under different GNN architectures.